You are considering a new product launch. The project will cost $860,000, have a 4-year life, and have no salvage value; depreciation is straight-line to zero. Sales are projected at 520 units per year; price per unit will be $18,800, variable cost per unit will be $15,500, and fixed costs will be $890,000 per year. The required return on the project is 13 percent, and the relevant tax rate is 24 percent. a. The unit sales, variable cost, and fixed cost projections given above are probably accurate to within ±10 percent. What are the upper and lower bounds for these projections? What is the base-case NPV? What are the best-case and worst-case scenarios? (A negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your NPV answers to 2 decimal places, e.g., 32.16.) b. Calculate the sensitivity of your base-case NPV to changes in fixed costs. (A negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.) c. What is the accounting break-even level of output for this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1
a] Base Case Upper Bound Lower Bound
Units per year 520 572 468
Variable cost per unit $                  15,500 $               17,050 $              13,950
Fixed costs $              8,90,000 $           9,79,000 $          8,01,000
BASE CASH NPV:
Sales = 520*18800 = $            97,76,000
Variable costs = 520*15500 = $            80,60,000
Fixed costs $              8,90,000
Depreciation = 860000/4 = $              2,15,000
NOI $              6,11,000
Tax at 24% $              1,46,640
NOPAT $              4,64,360
Add: Depreciation $              2,15,000
OCF $              6,79,360
PV of OCF = 679360*(1.13^4-1)/(0.13*1.13^4) = $            20,20,737
Less: Initial cost $              8,60,000
NPV $            11,60,737
BEST CASE NPV:
Sales = 572*18800 = $        1,07,53,600
Variable costs =572*13950 = $            79,79,400
Fixed costs $              8,01,000
Depreciation = 860000/4 = $              2,15,000
NOI $            17,58,200
Tax at 24% $              4,21,968
NOPAT $            13,36,232
Add: Depreciation $              2,15,000
OCF $            15,51,232
PV of OCF =1551232*(1.13^4-1)/(0.13*1.13^4) = $            46,14,095
Less: Initial cost $              8,60,000
NPV $            37,54,095
WORST CASE NPV:
Sales = 468*18800 = $            87,98,400
Variable costs = 468*17050 = $            79,79,400
Fixed costs $              9,79,000
Depreciation = 860000/4 = $              2,15,000
NOI $            -3,75,000
Tax at 24% $                -90,000
NOPAT $            -2,85,000
Add: Depreciation $              2,15,000
OCF $                -70,000
PV of OCF =70000*(1.13^4-1)/(0.13*1.13^4) = $            -2,08,213
Less: Initial cost $              8,60,000
NPV $          -10,68,213
b] Sensitivity of NPV to fixed costs = Change in NPV/Change in fixed costs = (3754095-1160737)/(810000-890000) = $                  -32.42
It means that for $1 change in fixed costs the NPV
will change in the opposite direction by $32.42.
A decrease of $89000 in fixed costs will increase
NPV by 89000*32.42 = $            28,85,380
c] Accounting break even level of output = Fixed costs/CM per unit = (890000+215000)/(18800-15500) = 335 Units